OpenStudy (anonymous):

Find the real solutions of the equation. 4(x+1)^2 + 14 (x+1) + 6 =0

jigglypuff314 (jigglypuff314):

if the problem was just 4x^2 + 14x + 6 = 0 what would you do?

OpenStudy (anonymous):

Factor?

jigglypuff314 (jigglypuff314):

first is to factor out the GCF then factor :)

OpenStudy (anonymous):

Which would be 2?

jigglypuff314 (jigglypuff314):

yes :)

OpenStudy (anonymous):

Ok, so 2(2+7+6) This doesn't look right to me. Can you please show me?

jigglypuff314 (jigglypuff314):

4x^2 + 14x + 6 = 0 factor out a 2 2(2x^2 + 7x + 3) = 0 then factor 2(2x+3)(x+2) = 0 understand this so far?

OpenStudy (anonymous):

Yes

jigglypuff314 (jigglypuff314):

so instead of x it was (x+1) so you would be getting 2 (2x +3) (x +2) = 0 2*(2(x+1) + 3)*((x+1) + 2) = 0 see?

OpenStudy (anonymous):

Yes, I think so

jigglypuff314 (jigglypuff314):

then 2*(2(x+1) + 3)*((x+1) + 2) = 0 2 ( 2x + 2 + 3) ( x + 1 +2) = 0 2 ( 2x + 5 ) ( x + 3 ) = 0

OpenStudy (anonymous):

Where did you get the 5 from?

jigglypuff314 (jigglypuff314):

2(x+1) + 3 = 2(x) + 2(1) + 3 = 2x + 2 + 3 = 2x + 5

OpenStudy (anonymous):

ok

jigglypuff314 (jigglypuff314):

so solve for x for 2x + 5 = 0 and x + 3 = 0

OpenStudy (anonymous):

I am studying for my midterm review, and on the answer sheet it says the answer is (-3/2,-4)

jigglypuff314 (jigglypuff314):

hmm, oh, sorry :/ I made a mistake 2(4(x+1)^2 +7(x+1) + 3) = 0 factors to 2(4(x+1) + 3)((x+1)+1) = 0 but I'm not sure how that is the answer :/ hmmm

OpenStudy (anonymous):

I am so confused!

jigglypuff314 (jigglypuff314):

sorry, would you like me to tag some people to help you?

OpenStudy (anonymous):

Yes if you could please.

jigglypuff314 (jigglypuff314):

@SolomonZelman @shamil98 @ehuman If you guys aren't too busy

OpenStudy (solomonzelman):

Let me think, hold on...

OpenStudy (solomonzelman):

-7

OpenStudy (solomonzelman):

@hilynn3, more help?

jigglypuff314 (jigglypuff314):

how did you get -7?

OpenStudy (solomonzelman):

I did it in my head, I'll explain....

OpenStudy (solomonzelman):

Let x+1=a OK?

OpenStudy (solomonzelman):

@60jigglypuff314, do you know how to get to it, or not? b/c I can't stay here unreplied for ever.

OpenStudy (solomonzelman):

@jigglypuff314

jigglypuff314 (jigglypuff314):

if you let (x+1) = x in 4(x+1)^2 + 14(x+1) + 6 = 0 then 4x^2 + 14x + 6 = 0 which you can factor to (4x+2)(x+3) = 0 then factor out a 2 2(2x+1)(x+3) = 0 then solve for the x's 2x+1 = 0 so 2x = -1 so x = -1/2 x + 3 = 0 so x = -3 and x = (x+1) so (x+1) = -1/2 so x = -1/2 - 1 = -3/2 (x+1) = -3 so x = -3 -1 = -4 so x = -3/2, -4

OpenStudy (solomonzelman):

4(x+1)^2 + 14 (x+1) + 6 =0 The answer is not -7. let (x+1) =a just to make it a little easier, so sub a for (x+1) 4a^2+14a+6=0 divide by 2 2a^2+7a+6=0 Can you guys solve it from there?

OpenStudy (solomonzelman):

Solve for a, and x will be EZ!

OpenStudy (anonymous):